# Conflicts in Texts: Data, Implications and Challenges

### v1

@@ 3.1 Origins
The subjective nature of human judgments bring noise and disagreements to their annotated data Kahneman (2021). Such annotation disagreements in NLP arise from multiple sources, including linguistic ambiguity, annotator backgrounds, task design, and dataset curation practices. Uma et al. (2021) provide a comprehensive survey of disagreement across NLP and computer vision tasks, highlighting subjective ambiguity and annotator diversity as key factors of disagreements. Sandri et al. (2023) categorize disagreements in offensive language detection, showing that some stem from inherent ambiguity while others result from annotation errors or lack of context. Their findings imply that not all disagreements are equal – some signal hard-to-classify content, whereas others indicate correctable annotation issues. Jiang and de Marneffe (2022) similarly classify NLI disagreements into three broad categories—linguistic uncertainty, annotator bias, and task design issues. They show that both linguistic uncertainty and annotator bias contribute substantially to label variability, and a significant portion of “disagreement noise” in NLI is systematic and predictable (e.g. stemming from specific ambiguity types or annotator profiles).
Task design also significantly influences annotation disagreements. Dsouza and Kovatchev (2025) find that labels disagreement in Reinforcement Learning from Human Feedback (RLHF) largely depend on annotator selection and task formulation, while Yung and Demberg (2025) show that free-choice annotation methods, where annotators select any suitable connective, yield less diversity (often converging on common labels) than forced-choice approaches, where annotators choose from predefined options, in discourse relation labeling.
On the other hand, demographic and ideological factors also shape disagreement. Pavlick and Kwiatkowski (2019) suggest that many NLI disagreements are not errors but reflect genuine ambiguities in language and individual variations in background knowledge. Sap et al. (2022) demonstrate that annotators’ personal beliefs and identities affect their perception of toxicity, and Wan et al. (2023) find that demographic data significantly enhances the prediction of annotation disagreements.
@@
Ethical and societal biases are also present in human-annotated texts. Biases introduced during annotation, whether related to race, gender, or geography, can significantly skew model predictions and decision-making processes Buolamwini and Gebru (2018). For instance, studies have shown that popular NLP datasets have a severe Western-centric skew Faisal et al. (2022). This Western dominance in training/evaluation data means models optimized on these benchmarks assume a Western context by default. A model might perform well on answering questions about New York or London (since those appear often in the data), but fail on questions about Nairobi or Manila simply due to lack of exposure Faisal et al. (2022).
Sap et al. (2022) show that annotators’ ideological and racial attitudes affect their judgments of toxicity, with conservative annotators less likely to flag anti-Black slurs as toxic but more likely to misclassify African American English (AAE) as offensive. Similarly, Thorn Jakobsen et al. (2022) examine how annotation guidelines interact with annotator demographics, finding that different task formulations can either exacerbate or mitigate biases. They show that even well-designed guidelines can elicit systematically different responses from distinct demographic groups, underscoring the importance of inclusive task framing to reduce disparities in annotations. Pei and Jurgens (2023), on the other hand, introduce POPQUORN, a dataset explicitly designed to measure the impact of annotator demographics across multiple NLP tasks. Their large-scale analysis confirms that annotator background, such as age, gender, race, and education, accounts for substantial variance in annotations.
@@ 3.1.1 Implications and Mitigation
Early research has highlighted the impact of annotator disagreements on data quality and model performance Artstein and Poesio (2008); Pustejovsky and Stubbs (2012); Plank et al. (2014). Pavlick and Kwiatkowski (2019) find that standard NLI models’ uncertainty did not reflect the true uncertainty found in human opinions. This mismatch suggests that when disagreements are ignored, models become overconfident on contentious cases. In addition, Anand et al. (2023) observe that a classifier provided only with single “gold” labels tends to be less confident and less accurate on examples where annotators widely disagreed on. Such models may treat these instances as “hard to learn” or mislabeled. Furthermore, Sap et al. (2019) reveal how annotator biases can lead to racist outcomes in hate speech classifiers. They uncover a spurious correlation: tweets written in African American English (AAE) are often rated as more toxic by annotators, even when they are not hate speech. Models trained on such data inherit this bias, falsely flagging content by Black authors as offensive at disproportionately high rates.
In terms of data collection, previous work has explored the strategy of acquiring multiple labels for each data item from various annotators to enhance data quality. They develop a probabilistic model to assess the true label of an item by considering the varying expertise of annotators and the possibility of label noise Sheng et al. (2008). Mostafazadeh Davani et al. (2022) examine strategies to train NLP models without discarding annotator disagreements. They propose a multi-annotator modeling approach: a multi-task neural network that learns to predict each individual annotator’s label as a separate output while sharing a common representation. Similarly, different studies have shown that models which incorporate annotator disagreement as "soft" labels (i.e. full label distribution) for training outperform those trained on aggregated single labels Uma et al. (2021); Fornaciari et al. (2021)
@@ 4.1.1 Origins
A common type of knowledge conflict arises when a model’s prompt (contextual knowledge) contradicts what the model has learned and stored in its parameters (parametric knowledge) Longpre et al. (2021); Chen et al. (2022a). One prevalent cause of such conflicts is the presence of updated information Chen et al. (2021a); Lazaridou et al. (2021); Luu et al. (2022), where newly available knowledge contradicts models’ previously learned knowledge.
@@
Different models may exhibit conflicts in their knowledge bases. Cohen et al. (2023) investigate how different LLMs can be used to fact-check each other to reveal inconsistencies that imply factually incorrect claims. On the other hand, Zhu et al. (2024) address the issue of inconsistencies between the visual and language components in Large Vision-Language Models (LVLMs). These conflicts arise due to the separate training processes and distinct datasets used for each modality, leading to discrepancies in the knowledge they capture. Even the same model may have internal knowledge conflicts. Zhao et al. (2024) identify intra-model contradictions by paraphrasing the same query multiple times and find answer divergence across different LLMs.
@@ 4.1.2 Implications and Mitigation
Various approaches have been proposed to mitigate the consequences of such knowledge conflicts. Longpre et al. (2021) propose a simple method that augments the training set with training examples modified by corpus substitution to mitigate memorization, Chen et al. (2022a) present a calibrator that abstains from predicting on instances with conflicting evidence, and Wang et al. (2024) propose a new instruction-based approach that augment LLMs to first identify knowledge conflicts, then pinpoint conflicting information segments, and lastly provide distinct answers in conflicting scenarios.
@@ 4.2.1 Origins
Factual hallucinations occur when a model’s output conflicts with real-world facts. TruthfulQA Lin et al. (2022) introduces an adversarial question-answering benchmark, revealing that even the best model at that time (GPT-3) was truthful on only 58% of questions, compared to 94% for humans. Pagnoni et al. (2021) construct the FRANK dataset, which annotates factual errors in summarization, identifying strengths and weaknesses of various metrics. Similarly, Honovich et al. (2021) extend QAGS to knowledge-grounded dialogue using question generation and entailment modeling to assess factual consistency. To further evaluate LLMs’ factual knowledge and reasoning, Hu et al. (2023) introduce Pinocchio, a benchmark comprising 20,713 multiple-choice questions across various domains, timelines, and languages, and assess models on fact composition, temporal reasoning, and adversarial robustness, revealing that LLMs often struggle with factual consistency and are prone to spurious correlations. Mallen et al. (2023) also find that language models struggle with less popular factual knowledge, and that retrieval augmentation helps significantly in these cases.
@@
Contextual hallucinations occur when generated text contradicts the given input context, such as in summarization, translation, and generation tasks. Maynez et al. (2020) find that summarization models frequently generate content unfaithful to input documents, with 64% of summaries containing unsupported information. In machine translation, Raunak et al. (2021) analyze hallucinations caused by source perturbations and training noise, and find that slight modifications to input data could trigger off-topic translations. Similarly, Dale et al. (2023) introduce HalOmi, a multilingual benchmark for hallucination and omission detection in machine translation, showing that prior hallucination detectors often fail across different language pairs. In generation tasks, Liu et al. (2021b) propose a novel token-level, reference-free hallucination detection task and dataset (HADES) for free-form text generation, and Niu et al. (2024) introduce RAGTruth, a comprehensive corpus designed for analyzing word-level hallucinations across various domains and tasks within standard Retrieval-Augmented Generation (RAG) frameworks.
@@ 4.2.2 Implications and Mitigation
5 Open Challenges
@@
Conflicts in training data and retrieved contexts can lead to unexpected consequences in NLP models, affecting their reliability and trustworthiness Pan et al. (2023); Liu et al. (2024). Studies show that knowledge conflicts within retrieved contexts severely degrade LLM performance in retrieval-augmented generation (RAG) Chen et al. (2022a); Pham et al. (2024); Liu et al. (2024); Li et al. (2024b), and QA models are highly susceptible to misinformation, where even minimal evidence contamination can lead to incorrect predictions Pan et al. (2023). Also, LLMs exhibit confirmation bias, where models prefer external evidence that aligns with their parametric knowledge, even when conflicting information is present Xie et al. (2024). This bias can cause models to reinforce incorrect or outdated information instead of updating their knowledge based on newly retrieved sources. We highlight the need of developing conflict-aware NLP systems that (1) assess and resolve contradictions in retrieved contexts, (2) mitigate confirmation bias by designing models that reason over conflicting evidence, and (3) integrate fact-checking mechanisms that evaluate source credibility and detect misinformation before generating responses.
Towards Robust and Fair NLP Models
Beyond technical challenges in handling knowledge conflicts, systemic biases in training data and annotation processes impact the fairness of NLP models. The distribution of web data skews LLM outputs toward Western perspectives, disproportionately reflecting Western English-speaking values Ramaswamy et al. (2023); Mihalcea et al. (2024); Tao et al. (2024); Naous et al. (2024). Furthermore, LLMs provide inconsistent answers across languages, revealing contradictions in how they encode cultural perspectives Li et al. (2024a); AlKhamissi et al. (2024). Biases also emerge from annotation processes, where demographic and ideological factors influence how data is labeled Sap et al. (2019). Future directions to create robust and fair NLP models include (1) building demographically diverse and geographically representative training datasets, (2) enhancing consistency of models by aligning cross-lingual knowledge representations, and (3) improving annotation frameworks to account for demographic and ideological biases among annotators.
@@
Conflicting information is present both in the data that models rely on and in their generated outputs. While we strive to account for all potential conflict scenarios, some cases may inevitably be overlooked. Additionally, due to space constraints, we do not provide an exhaustive discussion of the literature on each specific type of conflict. Instead, we adopt a broader perspective, examining various types of conflicts to identify connections, patterns, and common challenges.

### v2

@@ 3 Conflicts in Human-Annotated Texts
Conflicts in human-annotated texts largely arise from two sources: annotation disagreements and societal or ethical biases. Disagreements stem from linguistic ambiguity, annotator backgrounds, and task design, while biases reflect systematic demographic or ideological influences that can skew labeling in consistent ways. Though conceptually distinct, these sources often interact—biases may amplify disagreement or entrench disparities. Differentiating between them is essential for understanding annotation-related conflicts and for developing more reliable and equitable NLP datasets.
@@ 3.1 Origins
The subjective nature of human judgment introduces variability and disagreement into annotated data Kahneman (2021). In NLP, such disagreements arise from linguistic ambiguity, annotator backgrounds, task design, and dataset curation practices. Uma et al. (2021) survey disagreements across NLP and vision tasks, identifying subjective ambiguity and annotator diversity as key contributors. Sandri et al. (2023) classify disagreements in offensive language detection as stemming from inherent ambiguity, annotation errors, or contextual gaps, highlighting that some disagreements reflect hard-to-classify content, while others indicate correctable issues. Similarly, Jiang and de Marneffe (2022) categorize NLI disagreements into linguistic uncertainty, annotator bias, and task design, showing that much of the observed noise is systematic and predictable.
Task formulation also plays a critical role. Dsouza and Kovatchev (2025) find that label disagreement in reinforcement learning from human feedback (RLHF) is shaped by annotator selection and task phrasing. Demographic and ideological factors further influence disagreements. Pavlick and Kwiatkowski (2019) argue that many NLI disagreements reflect genuine linguistic ambiguity and individual variation rather than annotation error. Sap et al. (2022) demonstrate that annotators’ personal beliefs and identities affect toxicity judgments, while Wan et al. (2023) show that demographic features significantly improve disagreement prediction.
@@
Human-annotated texts also encode societal biases related to race, gender, and geography, which can significantly skew model predictions and downstream decisions Buolamwini and Gebru (2018). Sap et al. (2022) show that annotators’ ideological and racial identities influence toxicity judgments, with conservative annotators less likely to flag anti-Black slurs and more likely to misclassify African American English (AAE) as offensive. Thorn Jakobsen et al. (2022) examine how annotation guidelines interact with annotator demographics, demonstrating that even well-designed tasks can elicit systematically different responses across groups, highlighting the need for inclusive task design. Pei and Jurgens (2023) introduce POPQUORN, a dataset designed to assess demographic effects on annotation across NLP tasks, and find that annotator attributes—such as age, gender, race, and education—account for substantial variance in labeling behavior.
@@ 3.1.1 Implications and Mitigation
Implications
Early research has underscored the impact of annotator disagreement on data quality and model performance Artstein and Poesio (2008); Pustejovsky and Stubbs (2012); Plank et al. (2014). Pavlick and Kwiatkowski (2019) show that standard NLI models fail to capture the true uncertainty present in human judgments, leading to overconfidence on contentious examples. Similarly, Anand et al. (2024) find that models trained on single “gold” labels perform poorly and exhibit lower confidence on high-disagreement instances, often treating them as mislabeled or hard to learn. Sap et al. (2019) demonstrate how annotator bias can yield discriminatory outcomes: tweets in African American English (AAE) are frequently misclassified as toxic, a bias inherited by models that disproportionately flag content from Black authors. Additionally, many widely used NLP datasets exhibit a strong Western-centric skew Faisal et al. (2022), causing models to generalize poorly to underrepresented regions—for example, excelling on questions about New York or London, but failing on Nairobi or Manila due to lack of exposure.
Mitigation
Prior work has explored collecting multiple labels per data item to capture annotation variability and improve data quality. Probabilistic models have been developed to infer true labels by accounting for annotator expertise and label noise Sheng et al. (2008). Mostafazadeh Davani et al. (2022) propose a multi-task neural network that models each annotator’s labels individually while sharing a common representation, preserving disagreement in training. Similarly, studies show that models trained on soft labels—i.e., full label distributions reflecting annotator disagreement—consistently outperform those trained on aggregated labels Uma et al. (2021); Fornaciari et al. (2021).
@@ 4 Conflicts during Model Interactions
Conflicts during model interactions primarily manifest as knowledge conflicts and hallucinations, each posing distinct challenges. Knowledge conflicts occur when a model’s parametric memory contradicts contextual input or when inconsistencies arise across models, whereas hallucinations occur when outputs deviate from real-world facts or the given input. Differentiating these two types of conflict clarifies their underlying causes and helps guide targeted mitigation strategies.
@@ 4.1.1 Origins
A common type of knowledge conflict arises when a model’s prompt (contextual knowledge) contradicts what the model has learned and stored in its parameters (parametric knowledge) Longpre et al. (2021); Chen et al. (2022a). One prevalent cause of such conflicts is the presence of updated information Chen et al. (2021); Lazaridou et al. (2021); Luu et al. (2022), where newly available knowledge contradicts models’ previously learned knowledge.
@@
Conflicts may also arise across or within model knowledge bases. Cohen et al. (2023) explore how different LLMs encode different knowledge and can be used to fact-check one another, uncovering inconsistencies indicative of factual errors. Zhu et al. (2024) examine cross-modality conflicts in vision-language models, attributing discrepancies between visual and textual components to separate training regimes and distinct data sources. Even within a single model, contradictions can emerge: Zhao et al. (2024) detect intra-model inconsistencies by paraphrasing queries and observing divergent answers across prompts.
@@ 4.1.2 Implications and Mitigation
Mitigation
Several approaches have been proposed to mitigate the impact of knowledge conflicts. Longpre et al. (2021) reduce memorization by augmenting training data through corpus substitution. Chen et al. (2022a) introduce a calibrator that abstains from prediction when conflicting evidence is detected. More recently, Wang et al. (2024) propose an instruction-based framework that enables LLMs to identify conflicts, localize conflicting segments, and generate distinct responses for conflicting scenarios.
@@ 4.2.1 Origins
Factual hallucinations arise when a model’s output contradicts real-world facts. Lin et al. (2022) present TruthfulQA, an adversarial QA benchmark, and show that even top-performing models like GPT-3 were truthful on only 58% of questions, compared to 94% for humans. Pagnoni et al. (2021) construct FRANK, a dataset for identifying factual errors in summarization, while Honovich et al. (2021) extend QAGS to dialogue by leveraging question generation and entailment for factual consistency evaluation. To assess factual knowledge and reasoning in LLMs, Hu et al. (2024) introduce Pinocchio, a large benchmark covering multiple domains, timelines, and languages, revealing challenges in composition, temporal reasoning, and robustness. Mallen et al. (2023) further find that models struggle with less common factual knowledge, with retrieval augmentation significantly improving performance in such cases.
@@
Contextual hallucinations occur when generated text contradicts the given input context, such as in summarization, translation, and generation tasks. Maynez et al. (2020) find that summarization models frequently generate content unfaithful to input documents, with 64% of summaries containing unsupported information. In machine translation, Raunak et al. (2021) analyze hallucinations caused by source perturbations and training noise, and find that slight modifications to input data could trigger off-topic translations. Similarly, Dale et al. (2023) introduce HalOmi, a multilingual benchmark for hallucination and omission detection in machine translation, showing that prior hallucination detectors often fail across different language pairs. In generation tasks, Liu et al. (2022a) propose a novel token-level, reference-free hallucination detection task and dataset (HADES) for free-form text generation, and Niu et al. (2024) introduce RAGTruth, a comprehensive corpus designed for analyzing word-level hallucinations across various domains and tasks within standard Retrieval-Augmented Generation (RAG) frameworks.
@@ 4.2.2 Implications and Mitigation
Implications
Hallucinations threaten trust, safety, and the integrity of AI-powered workflows. Hallucinated outputs can rapidly spread false information. For instance, in 2023, an AI-generated image purporting to show an explosion near the Pentagon went viral, briefly causing public panic and even a stock market dip before being debunked Sun et al. (2024). Hallucinations directly degrade the performance of downstream applications like abstractive summarization. Studies have found that a large portion of generated summaries contain unsupported facts, misleading readers and propagating misinformation in news and scientific dissemination Kryscinski et al. (2020).
Mitigation
@@
5 Connections, Challenges and Directions
Given the significance and impact of conflicts in NLP, we advocate for increased attention to the development of conflict-aware and robust AI systems. In this section, we highlight specific challenges by connecting different types of conflicts and propose concrete research directions to address them.
Culturally Robust LLMs
Among the challenges outlined in this survey, the development of culturally robust LLMs remains particularly underexplored. Cultural conflicts emerge both in naturally occurring web data and human-annotated datasets, where Western-centric distributions dominate. Prior studies reveal that LLMs often reflect the values and perspectives of Western, English-speaking populations Ramaswamy et al. (2023); Mihalcea et al. (2025); Tao et al. (2024); Naous et al. (2024), with misalignments especially pronounced for underrepresented personas and culturally sensitive topics Al Kuwatly et al. (2020). Additionally, LLMs exhibit inconsistent behavior across languages Li et al. (2024a); AlKhamissi et al. (2024); Eloundou et al. (2025), revealing internal cultural conflicts. These issues are rooted in the data: both the pre-train data and benchmark datasets commonly exhibit Western-centric biases Mihalcea et al. (2025); Faisal et al. (2022), causing models to default to Western contexts and perform poorly on less-represented regions and cultures.
However, to the best of our knowledge, no effective methodology has yet been proposed to address this issue. With the emergence of culturally distinct LLMs—such as Qwen, trained largely on Chinese data Bai et al. (2023), and Vikhr, trained on Russian data Nikolich et al. (2025)—a promising direction is model fusion across culturally diverse models to achieve greater cultural balance Wan et al. (2024b); Jiang et al. (2023a). Furthermore, advances in culture-specific LLMs and synthetic data generation offer the potential to curate more culturally representative training and evaluation datasets beyond Western-centric narratives, supporting the development of culturally robust LLMs.
@@ 5 Connections, Challenges and Directions
As outlined in this survey, various types of conflicts can arise in a model’s input, each requiring different handling depending on the task. We argue that downstream applications should not treat all conflicts uniformly; rather, responses should be tailored to the conflict type. For instance, conflicts due to ambiguity should elicit clarification questions, factual contradictions should trigger reasoning over evidence, and opinion-based disagreements should induce balanced, multi-perspective responses. Realizing such capabilities requires models to be aware of the potential conflicts and classify them according to a systematic taxonomy. Yet, current research lacks frameworks to distinguish and operationalize these conflict types. Our proposed taxonomy offers a foundational step toward enabling conflict-aware systems that can recognize, interpret, and appropriately address diverse conflicts in downstream applications.
6 Conclusion
We present a unified view of conflicting information in NLP, organizing the landscape into conflicts originating from (i) natural texts on the web, (ii) human annotations, and (iii) model interactions. This taxonomy connects lines of work that are often studied in isolation and clarifies how conflicts arise, what they imply for reliability, and how current methods aim to mitigate them. Our synthesis argues that conflict awareness should guide the full pipeline: data collection that preserves disagreement and multiple perspectives, models that detect and categorize conflicts and respond with clarification, reasoning, or balanced presentation, and evaluations that measure calibration under disagreement, robustness to contradictory evidence, and cultural coverage across languages and regions.
@@ 6 Conclusion
Conflicting information is present both in the data that models rely on and in their generated outputs. While we strive to account for all potential conflict scenarios, some cases may inevitably be overlooked. Additionally, due to space constraints, we cannot provide an exhaustive discussion of the literature on each specific type of conflict. Instead, we adopt a broader perspective, examining various types of conflicts to identify connections, challenges, and future directions.
Acknowledgment
This research is based upon work supported in part by the Office of the Director of National Intelligence (ODNI), Intelligence Advanced Research Projects Activity (IARPA), via 2022-22072200003. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of ODNI, IARPA, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein. This work is also supported by NSF grant #IIS-2135581.
Appendix A Summary Tables
In the summary tables, dataset covers prior work that proposed datasets and benchmark, method covers work that focus on mitigation strategies, and analysis presents work that aim at providing insights through experiments.
Table 1: Datasets, methods, and analysis for conflicts in natural texts
Conflict Type Sub-type Category Work
Factual Ambiguity Dataset SituatedQA Zhang and Choi (2021)
AmbigQA Min et al. (2020)
Time-sensitive QA Chen et al. (2021)
StreamingQA Liska et al. (2022)
Real-time QA Kasai et al. (2023)
Method Disambiguate then answer Cole et al. (2023)
Time-aware LM Dhingra et al. (2022)
Contradictory Evidence Dataset QACC Liu et al. (2024)
Contra-Doc Li et al. (2024b)
WhoQA Pham et al. (2024)
Machine-generated Pan et al. (2023)
Machine-generated Wan et al. (2024a)
Machine-generated Hong et al. (2024)
Machine-generated Jiayang et al. (2024)
Entity-substitution Chen et al. (2022a)
Rule-based Kazemi et al. (2023)
Method Finetuned Calibrator Chen et al. (2022a)
Finetuned w/ Explanation Liu et al. (2024)
Finetuned discriminator Hong et al. (2024)
Analysis Confirmation bias Xie et al. (2024)
Opinion Perspectives Dataset PERSPECTRUM Chen et al. (2019)
Multi-OpEd Liu et al. (2021)
NeuS Lee et al. (2022)
ConflictingQA Wan et al. (2024a)
Reddit Plepi et al. (2024)
Method Multi-task learning Liu et al. (2021)
Opinion summarization Jiang et al. (2023b)
Tailored generation Plepi et al. (2024)
Framing Bias Dataset MFC Card et al. (2015)
GVFC Liu et al. (2019)
BASIL Fan et al. (2019)
Method Multifaceted analysis Milbauer et al. (2021)
Pre-training Liu et al. (2022b)
Disentanglement Chen et al. (2023)
Multi-task learning Lee et al. (2022)
Analysis Sentence-level Lei et al. (2022)
Table 2: Datasets, methods, and analysis for conflicts in human-annotated texts
Conflict Type Sub-type Category Work
Human-Annotated Disagreement Dataset Twitter Sandri et al. (2023)
RLHF Dsouza and Kovatchev (2025)
DiscoGeM Yung and Demberg (2025)
NLI Pavlick and Kwiatkowski (2019)
Method Probabilistic model Sheng et al. (2008)
Multi-task Mostafazadeh Davani et al. (2022)
Soft labels Uma et al. (2021)
Soft labels Fornaciari et al. (2021)
Analysis Survey Uma et al. (2021)
Survey Klie et al. (2023)
Offensive language Sandri et al. (2023)
NLI Jiang and de Marneffe (2022)
Task design Dsouza and Kovatchev (2025)
Free choice Yung and Demberg (2025)
Personal belief Sap et al. (2022)
Demographic data Wan et al. (2023)
Biases Dataset Gender Buolamwini and Gebru (2018)
Argument mining Thorn Jakobsen et al. (2022)
POPQUORN Pei and Jurgens (2023)
Analysis Western-centric Faisal et al. (2022)
Toxicity Wan et al. (2023)
Racist outcome Sap et al. (2019)
Table 3: Datasets, methods, and analysis for conflicts during model interactions
Conflict Type Sub-type Category Work
Knowledge Context vs. Memory Dataset Entity substitution Longpre et al. (2021)
Entity substitution Chen et al. (2022a)
Instruction-based Wang et al. (2024)
Misinformation injection Pan et al. (2023)
KRE Ying et al. (2024)
context-conflicting Tan et al. (2024)
Method Data Augmentation Longpre et al. (2021)
Abstention Chen et al. (2022a)
Instruction-based Wang et al. (2024)
Within & Across Analysis LM-vs-LM fact-checking Cohen et al. (2023)
Cross-modality Zhu et al. (2024)
Intra-model contradiction Zhao et al. (2024)
Hallucination Factual Dataset TruthfulQA Lin et al. (2022)
FRANK Pagnoni et al. (2021)
$q^{2}$ Honovich et al. (2021)
Pinocchio Hu et al. (2024)
MiniCheck Tang et al. (2024)
Method RAG Lewis et al. (2020)
RAG Shuster et al. (2021)
Knowledge graph Guan et al. (2024)
Disentanglement Neeman et al. (2023)
QA validation Kryscinski et al. (2020)
QA validation Wang et al. (2020)
Entailment-based Laban et al. (2022)
SelfCheckGPT Manakul et al. (2023)
Analysis Less popular entities Mallen et al. (2023)
Contextual Dataset HalOmi Dale et al. (2023)
HADES Liu et al. (2022a)
RAGTruth Niu et al. (2024)
Method Context-aware decoding Shi et al. (2024)
Long context Liu et al. (2025)
Context-DPO Bi et al. (2024)
CR-DPO Huang et al. (2025)
Analysis Summarization Maynez et al. (2020)
Translation Raunak et al. (2021)
Appendix B Survey Methodology
Our objective is to map the broad and diverse landscape of textual conflicts in NLP, identify central implications and gaps, and outline a research agenda. We surveyed research indexed in Google Scholar, the ACL Anthology, and OpenReview, covering peer-reviewed work in major ML and NLP venues such as ICLR and ACL, as well as recent manuscripts on arXiv. Rather than relying on predetermined keyword filters, we primarily used citation chaining: starting from influential surveys corresponding to the three sources of conflict considered in this work (natural web text, human annotations, and model interactions), we applied both backward and forward chaining using Google Scholar and venue indexes. Paper selection did not follow a rigid checklist. The two authors independently screened and included papers based on inclusion criteria that prioritized recency, citation impact, and perceived influence on the area, then reconciled disagreements through discussion. This methodology emphasizes coverage and structure over exhaustiveness and is intended to synthesize a rapidly evolving field while making the scope of inclusion explicit and surfacing open problems.